A painting measuring 12 inches by 18 inches is surrounded by a frame of uniform width. If the combined area of the painting and the frame is 432 square inches, determine the width of the frame.
Answer by checkley75(3666)

(12+2x)(18+2x)=432
216+36x+24x+4x^2=432
4x^2+60x+216-432=0
4x^2+60x-216=0
x^2+15x-54=0
(x-18)(x+3)=0
x-18=0
x=18 not an answer.
x-3=0
x=3 answer for the width of the frame.
proof
(12+6)(18+6)=432
18*24=432
432=432